OpenStudy (anonymous):

let an=na n-2,with a1=1and a2=2.then a4a3 equals A.12 B.24 C.2 D.7

OpenStudy (anonymous):

\[a_n=na_{n-2}\] with \[a_1=1,a_2=2\]?

OpenStudy (anonymous):

thats correct i just couldnt type it like that

OpenStudy (anonymous):

that is my guess in any case first you need \(a_3\) which is \[3\times a_{3-2}=3\times a_1=3\times 1=3\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then you need \(a_4\) which is \[4\times a_{4-2}=4a_2=4\times 2=8\]

OpenStudy (anonymous):

ok , can u help me solve the rest

OpenStudy (anonymous):

what is left to do?

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

you were asked for \(a_3\times a_4\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

we found \(a_3=3,a_4=8\) all that is left for you is to multiply

OpenStudy (anonymous):

so it would be 3x8=24

OpenStudy (anonymous):

@satellite73 can u help me with absolute value